The kinetic energy of 5.0 kg object is initially 35 j. It is then subject to a constant force so that after traveling 9.5 m, its speed is doubled. Find (a)its initial speed (b)its kinetic energy after traveling 9.5 m and (c)the constant force acting on it

1
Expert's answer
2021-12-07T20:12:41-0500

(a) "KE_1=mv^2/2\to v=\sqrt{2KE/m}=\sqrt{2\cdot35/5}=3.74\ (m/s)"


(b) "KE_2=m(2v)^2/2=4(mv^2/2)=4\cdot KE_1=4\cdot35=140\ (J)"


(c) "\Delta KE=W\to 140-35=F\cdot9.5\to F=11\ (N)"